A line is equally inclined at an angle ‘a’ to the x and z axis.
If it makes angle ‘B’ with the y axis and (sinB)^2=3*(sina)^2,find a

Grade:12

1 Answers

Aditya Gupta
5 years ago
2cos^2a+cos^2b=1
or 2cos^2a= sin^2b
also  (sinB)^2=3*(sina)^2
so 2cos^2a= sin^2b= 3*(sina)^2
or 2=5*(sina)^2
root(2/5)= sina
a= arcsin(root(2/5))